OpenStudy (anonymous):

Use the elimination method to solve the system for x and y. 16x+7y=25 4+8x-2y=0 Select one: a. (3, 1/4) b. (1/4, -3) c. (-1/4, -3) d. (1/4, 3)

OpenStudy (anonymous):

1: 16x+7y=25 (multiply by 2) you get: 32x+14y=50 2: 4+8x-2y=0 (rearranging you get 8x-2y=-4) multiply by 7: you get 56x-14y=-28 now add the 1 and 2, you get: 32x+14y=50 + 56x-14y=-28 _________________ 88x=22 x=22/88 x=1/4 substitute into x=1/4 into 1: 16(1/4)+7y=25 4+7y=25 7y=25-4 7y=21 y=21/7 y=3 Therefore answer is d. (1/4,3)

OpenStudy (anonymous):

thanks(: